OpenStudy (anonymous):

An artillery shell is fired at an angle of 62.7 above the horizontal ground with an initial speed of 1520 m/s. The acceleration of gravity is 9.8 m/s2 . Find the total time of flight of the shell, neglecting air resistance. Answer in units of min

OpenStudy (anonymous):

vertical is 1520sin62.7 = 1350.69819 horizontal is 1520cos62.7 = 697.14732 s=ut +1/2 at^2 therefore 1350.698t +4.9t^2 t=277.77 seconds... and then divide 697.1432/277.77 = part b Can someone tell me if this is the correct answer, or if not what I am doing wrong.

OpenStudy (anonymous):

isnt this a math question?

OpenStudy (anonymous):

nope... university physics

OpenStudy (anonymous):

ok did not see word gravity

OpenStudy (anonymous):

no help ? :(

OpenStudy (anonymous):

looking im trying to figure out

OpenStudy (anonymous):

ok. thankyou

OpenStudy (anonymous):

lol i can only see the first part of what you typed the rest i got lost

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

it got too crowded up

OpenStudy (anonymous):

Assuming you plugged the values correctly into your calculator, that's correct. The procedure you went through is right. You need to convert from seconds to minutes, though. For part b, I assume you're looking for the total range of the projectile? In which case you need multiply the time of flight (in seconds) by the velocity.

OpenStudy (anonymous):

hey eashmore

OpenStudy (anonymous):

kyrstal. Didn't I answer this question for you already?

OpenStudy (anonymous):

it is nt right :(

OpenStudy (anonymous):

uhm

OpenStudy (anonymous):

You need to watch your units. Your answer should be in minutes, not seconds, right?

OpenStudy (anonymous):

im pulling it u one second.

OpenStudy (anonymous):

up*

OpenStudy (anonymous):

Well in your question it says minutes...

OpenStudy (anonymous):

are you good with vectors?

OpenStudy (anonymous):

it was right :/ just didnt read the whole problem.

OpenStudy (anonymous):

The work you did was correct. As Jemurray3 correctly pointed out, you need to convert seconds to minutes.

OpenStudy (anonymous):

can you help me with a math problem?